OpenStudy (anonymous):

True or false: If an angle is acute, then its complement must be greater than its supplement.

OpenStudy (anonymous):

I'm pretty sure it's false, but I don't know how to explain why it's false.

OpenStudy (anonymous):

False

OpenStudy (anonymous):

Yes it is..

OpenStudy (anonymous):

Ok, but how is it false?

OpenStudy (anonymous):

If it is acute, then its compliment is \(90-x\) And its Suppliment will be : \(180-x\) Now, \(0 \le x \le 90\), let us say x = 45, Then Compliment will be : \(90-45 = 45\) And its suppliment will be : \(180-45 = 135\)..

OpenStudy (anonymous):

Its because a supplementary angle has 180 degrees and complementary angle has to be 90 degrees so a complement angle cant be as big then the supplement angle. I think

OpenStudy (anonymous):

Obviously, Supplement \((135)\) is greater than Complement (\(45\))..

OpenStudy (anonymous):

Wow that helps a LOT!! Thanks!!!